# A large Deviation Principle for the Normalized Excursion of $\alpha$-Stable Lévy Processes without Negative Jump

### v2

@@ 2.2.  Basic results on the Skorokhod M1’ topology
In this section, we introduce the distance $\mathrm{dist}$ on $\mathbb{D}[0,1]$ inducing the Skorokhod M1’ topology considered in [2, 30], and study some of its key properties. We also refer to [31, Chapters 12 and 13] for a complete treatment of the Skorokhod M1 topology.
@@ 2.2.1.  The space $(\mathbb{D}[0,1],\mathrm{dist})$
This indeed defines a distance function that makes $(\mathbb{D}[0,1],\mathrm{dist})$ a Polish space [^1] Note, however, that the distance $\mathrm{dist}$ is not complete, see [31, Section 12.8].. These results are obtained exactly as in the context of the M1 topology [31, Theorems 12.3.1 and 12.8.1], with the slight modification that the convention taken in this reference is that $f(0-)=f(0)$ rather than $f(0-)=0$, and that $f$ is supposed to be continuous at 00 and $1$. The choice of convention that $f(0-)=0$ allows a sequence of functions that jump “right after time 0” to be possibly convergent in the M1’ topology. For example, one has $\mathrm{dist}(\mathbbm{1}_{[1/n,1]},\mathbbm{1}_{[0,1]})\to 0$ as $n\to\infty$, while the sequence $\mathbbm{1}_{[1/n,1]}$ is not convergent in the classical M1 topology, and in fact the M1’ topology is strictly weaker than the M1 topology. On the other hand, observe that $\mathbbm{1}_{[0,1/n]}$ is not convergent in $(\mathbb{D}[0,1],\mathrm{dist})$.
@@ 2.2.2.  The $M$-oscillation
This theorem can be found in Chapter 12 of Whitt [31], with the minor difference that our space of functions starts with an “initial jump” (recall that $f(0-)=0$ by our convention).
@@ 2.2.3.  An exponential tightness criterion
For 1., we follow and adapt the approach of Billingsley [4]. Let $D_{n}=\{k2^{-n},0\leq k\leq 2^{n}\}$ be the dyadic numbers of level $n$. Then for $\beta>0$ and $\lambda>0$, by Markov’s inequality,
@@
which by optimizing over $\lambda>0$ and taking a union bound yields
@@
Setting $A_{n}=\max\{M(f(k2^{-n},(k+1)2^{-n},(k+2)2^{-n})),0\leq k\leq 2^{n}-2\}$, this shows that $\mathbb{P}\left(A_{n}>2^{-n\beta}\right)\leq 2^{n}c\exp(-C^{\prime}(\alpha)\varepsilon^{-\alpha^{\prime}}2^{n(1-\alpha\beta)/(\alpha-1)})$. Next, let $f\in\mathbb{D}[0,1]$ and, for $I\subset[0,1]$, let
@@ 3.   Large deviations for the finite-dimensional marginal distributions
Let $\sigma=(t_{1},\ldots,t_{n})$, where $10<t_{1}<\dots<t_{n}<$ 1, be fixed. Under $\mathbb{P}$ the laws of $\varepsilon(\mathbbm{e}_{t_{1}},\dots,\mathbbm{e}_{t_{n}})$ satisfy a LDP in $\mathbb{R}^{n}$ with speed $\varepsilon^{-\alpha^{\prime}}$ and good rate function
@@ 3.1.  Estimates for transition densities
First, [27, Equation (14.35)] entails that for every $x\geq 0$, we have
$c_{1}\exp\left(-c_{\alpha}x^{\alpha^{\prime}}\right)\leq p_{1}(-x)\leq c_{2}(1+x^{\frac{2-\alpha}{2\alpha-2}})\exp\left(-c_{\alpha}x^{\alpha^{\prime}}\right)$ (3.1)
and [27, Equation (14.34)] entails that
@@
Next, by [23, Formula (3.20)], it holds that there exists a positive constant $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0})$ and for $0<a<b<+\infty$,
@@
These estimates will allow us to evaluate the densities involved in (2.3) in the large deviations regime. First we give an explicit formula for $p^{(0,\infty)}_{t}(x,y)$.
@@
Using the scaling properties of $p_{t}(x)$ and $q_{x}(t)$, we may deduce from 3.2 a bound on the error when we approximate $p^{(0,\infty)}_{t}(x,y)$ by $p_{t}(y-x)$, as follows.
@@
$\int_{0}^{t}q_{x}(s)p_{t-s}(y)\,\mathrm{d}s$ $\leq c_{2}\|p_{t/2}\|_{\infty}\int_{0}^{t/2}\frac{x\,\mathrm{d}s}{(t/2)^{\alpha}s^{1+1/\alpha}}(1+(x/s^{1/\alpha})^{\frac{2-\alpha}{2\alpha-2}})\exp\left(-c_{\alpha}(x/s^{1/\alpha})^{\alpha^{\prime}}\right)$
$+c_{2}\|p_{1}\|_{\infty}\frac{x}{(t/2)^{1+1/\alpha}}\left(1+\left(\frac{x}{(t/2)^{1/\alpha}}\right)^{\frac{2-\alpha}{2\alpha-2}}\right)\exp\left(-c_{\alpha}\left(\frac{x}{t^{1/\alpha}}\right)^{\alpha^{\prime}}\right)\int_{t/2}^{t}\frac{\,\mathrm{d}s}{(t-s)^{1/\alpha}}.$
@@
Proof of 3.1.
Since $J_{\sigma}$ is a good rate function on $\mathbb{R}^{n}$, [29, Lemma 5] shows that it suffices to prove that, for every open subset $G\subset\mathbb{R}^{n}$,
@@
$\Psi_{\varepsilon}(x_{1},\dots,x_{n})\coloneqq\prod_{i=1}^{n-1}p_{t_{i+1}-t_{i}}^{(0,\infty)}\Big{(}\frac{x_{i}}{\varepsilon},\frac{x_{i+1}}{\varepsilon}\Big{)}\times q_{x_{n}/\varepsilon}(1-t_{n}).$
@@
$p^{(0,\infty)}_{t_{i+1}-t_{i}}\left(\frac{x_{i}}{\varepsilon},\frac{x_{i+1}}{\varepsilon}\right)\geq c_{1}\exp\left(-\frac{c_{\alpha}}{\varepsilon^{\alpha^{\prime}}}\left(\frac{(b_{i}-a_{i+1})^{\alpha}}{t_{i+1}-t_{i}}\right)^{\frac{1}{\alpha-1}}\right)-C(\eta)\exp\left(-\frac{c_{\alpha}}{\varepsilon^{\alpha^{\prime}}}(1-\eta)\left(\frac{a_{i}^{\alpha}}{t_{i+1}-t_{i}}\right)^{\frac{1}{\alpha-1}}\right),$
@@ 4.   Exponential tightness for the normalized excursion
Now by [23, (3.20)], we have that $\bar{J}_{1}(x)\sim\bar{c}x^{-\alpha}$ as $x\to\infty$ for some finite constant $\bar{c}$, and the desired uniform upper bound follows. ∎
@@
$\mathbb{P}\left(\sup_{1-\delta<t\leq 1}\varepsilon\mathbbm{e}_{t}\geq\gamma,\varepsilon\mathbbm{e}_{1-\delta}<\gamma\right)\leq\alpha\Gamma\left(1-\frac{1}{\alpha}\right)\int_{0}^{\gamma/\varepsilon}\mathrm{d}x\,j_{1-\delta}(x)C(\eta,\gamma)\exp\bigg{(}-(1-\eta)c_{\alpha}\left(\frac{\gamma^{\alpha}}{\varepsilon^{\alpha}\delta}\right)^{\frac{1}{\alpha-1}}\bigg{)}.$
@@
Proof of 4.1.
Fix $\lambda>0$. By 4.3, for every $n\geq 1$, there exists a $\delta_{n}\in(0,1)$ such that for every $\varepsilon>0$ small enough
@@
For this choice of $\delta_{n}$, by 4.2 and 2.2, for every $n\geq 0$ there exists a compact set $K^{(n)}_{\lambda}$ of $\mathbb{D}[0,1]$ such that for every $\varepsilon>0$ small enough,
@@
5.   Proof of 1.2
In this section, we prove 1.2. We will do this by combining the exponential tightness with a weaker form of the LDP, as we now explain. The weak topology $\mathcal{W}$ on $\mathbb{D}[0,1]$ is the topology generated by the basis of neighborhoods of the form
@@
where $f\in\mathbb{D}[0,1]$, $\varepsilon_{1},\ldots,\varepsilon_{k}>0$, and $t_{1},\ldots,t_{k}$ are elements of $[0,1]$ that are continuity points of $f$. Here, by convention, 00 is a continuity point of $f$ if and only if $f(0)=0$, which is consistent with our convention that $f(0-)=0$. Clearly this defines a Hausdorff topology, since two different elements of $\mathbb{D}[0,1]$ necessarily differ at some common continuity point. It is easy to see that a sequence $(f_{n},n\geq 0)$ that converges to a limit $f$ in $(\mathbb{D}[0,1],\mathrm{dist})$ also converges pointwise at every continuity point of $f$, and, therefore, will eventually belong to any given basic neighborhood $N(f,t_{1},\ldots,t_{k},\varepsilon_{1},\ldots,\varepsilon_{k})$ for the topology $(\mathbb{D}[0,1],\mathcal{W})$. Since open sets in the metric space $(\mathbb{D}[0,1],\mathrm{dist})$ can be characterized sequentially, this implies that the weak topology is coarser than the topology of $(\mathbb{D}[0,1],\mathrm{dist})$. Therefore, by [9, Corollary 4.2.6], and by the exponential tightness established in 4.1, 1.2 will follow from the following statement.
@@
The remainder of this section is thus devoted to the proof of this proposition, which follows the approach of Lynch and Sethuraman [20] closely.
@@ 5.1.  Facts about the rate function
where, by convention, $x_{n+1}=0$ and $t_{n+1}=1$. We let $J_{\sigma}(x_{1},\ldots,x_{n})=\infty$ if one of the $x_{i}$’s is negative. To ease notation, for $f:[0,1]\to\mathbb{R}$, we let $I_{\mathbbm{e}}^{\sigma}(f)=J_{\sigma}(f(t_{1}),\ldots,f(t_{n}))$. By the Dawson-Gärtner theorem [9, Theorem 4.6.1], it follows from Proposition 3.1 that the laws of $(\varepsilon\mathbbm{e}_{t})_{t\in[0,1]}$ satisfy an LDP in $\mathbb{R}^{[0,1]}$ (with the product topology) as $\varepsilon\downarrow 0$, with speed $\varepsilon^{-\alpha^{\prime}}$ and good rate function
@@
We cannot immediately make use of this, since the domain of this rate function is not a space of càdlàg functions. However, let us prove some properties of the rate function $\widetilde{I}_{\mathbbm{e}}$ and, in particular, that its restriction to $\mathbb{D}[0,1]$ coincides with the rate function $I_{\mathbbm{e}}$ given in 1.2. To this end, we prove the following proposition.
@@
$M_{\mathbbm{e}}(f)\coloneqq\sup_{\sigma\in\mathfrak{S}}I_{\mathbbm{e}}^{\sigma}(f)<\infty.$
@@
This statement should be compared with [20, Theorem 3.2], where the proof uses a martingale argument. We provide another elementary proof here, based on the Lebesgue differentiation theorem instead. For convenience, let $\Lambda(x)=c_{\alpha}(x_{-})^{\alpha^{\prime}}$ for all $x\in\mathbb{R}$.
@@
By the Lebesgue decomposition theorem, we may write $f_{\uparrow}=f_{\uparrow\mathrm{ac}}+f_{\uparrow\mathrm{sing}}$, where $f_{\uparrow\mathrm{ac}}$ is an absolutely continuous function and $f_{\uparrow\mathrm{sing}}$ is such that $df_{\uparrow\mathrm{sing}}$ is singular with respect to the Lebesgue measure. By the Lebesgue differentiation theorem, for Lebesgue-almost every $t\in[0,1]$ we have
@@
Since the weak topology is not first-countable, we must initially use nets to characterise the lower-semicontinuity of $I_{\mathbbm{e}}$. We first need to show that if $(f_{\lambda})$ is a net that converges to $f$ in the weak topology, with $f_{\lambda}\geq 0$ and $f_{\lambda}(1)=f(1)=0$ for every $\lambda$, then $\liminf_{\lambda}\ I_{\mathbbm{e}}(f_{\lambda})\geq I_{\mathbbm{e}}(f)$. Let $\sigma=(t_{1},\ldots,t_{k})$ be a subdivision of continuity points of $f$, so that $f_{\lambda}(t_{i})$ converges to $f(t_{i})$ for $1\leq i\leq k$, and $I^{\sigma}_{\mathbbm{e}}(f_{\lambda})=J_{\sigma}(f_{\lambda}(t_{1}),\ldots,f_{\lambda}(t_{k}))$ converges to $I^{\sigma}_{\mathbbm{e}}(f)=J_{\sigma}(f(t_{1}),\ldots,f(t_{k}))$. Since $I_{\mathbbm{e}}(f_{\lambda})\geq I^{\sigma}_{\mathbbm{e}}(f_{\lambda})$ by 5.2, this implies that $\liminf_{\lambda}I_{\mathbbm{e}}(f_{\lambda})\geq I^{\sigma}_{\mathbbm{e}}(f)$. Applying 5.2 once again allows us to conclude that $I_{\mathbbm{e}}$ is a rate function on $(\mathbb{D}[0,1],\mathcal{W})$, and therefore also on $(\mathbb{D}[0,1],\mathrm{dist})$.
@@
The proof follows that of [20, Theorem 3.5] closely. Since we know from Proposition 5.2 that $I^{\sigma}_{\mathbbm{e}}(A)\leq I_{\mathbbm{e}}(A)$ for every $\sigma\in\mathfrak{S}$ and every set $A$, let us assume, for a contradiction, that $\sup_{\sigma\in\mathfrak{S}}I^{\sigma}_{\mathbbm{e}}(F)<c<I_{\mathbbm{e}}(F)$ for some constant $c\in(0,\infty)$. For every subdivision $\sigma=(t_{1},\ldots,t_{k})\in\mathfrak{S}$, we may find some element $f_{\sigma}\in\mathbb{D}[0,1]$ such that $I^{\sigma}_{\mathbbm{e}}(f_{\sigma})<c$. Let $\hat{f}_{\sigma}$ be the piecewise affine interpolation of the values of $f_{\sigma}$ at times $10<t_{1}<\ldots<t_{k}<$ 1, with $\hat{f}_{\sigma}(0)=\hat{f}_{\sigma}(1)=0$ and of course $\hat{f}_{\sigma}(t_{i})=f_{\sigma}(t_{i}),1\leq i\leq k$. Then, plainly,
@@
We now have all the tools needed to prove Proposition 5.1. The proof is split into two lemmas which follow [20, Theorems 4.1 and 4.2] closely.

### v3

@@ 2.2. Basic results on the Skorokhod M1’ topology
In this section, we introduce the distance $\mathrm{dist}$ on $\mathbb{D}[0,1]$ inducing the Skorokhod M1’ topology considered in Bazhba et al. (2020); Vysotsky (2021), and study some of its key properties. We also refer to (Whitt, 2002, Chapters 12 and 13) for a complete treatment of the Skorokhod M1 topology.
@@ 2.2.1. The space $(\mathbb{D}[0,1],\mathrm{dist})$
This indeed defines a distance function that makes $(\mathbb{D}[0,1],\mathrm{dist})$ a Polish space [^1] Note, however, that the distance $\mathrm{dist}$ is not complete; see (Whitt, 2002, Section 12.8).. These results are obtained exactly as in the context of the M1 topology (see (Whitt, 2002, Theorems 12.3.1 and 12.8.1)), with the slight modification that the convention taken in this reference is that $f(0-)=f(0)$ rather than $f(0-)=0$, and that $f$ is supposed to be continuous at 00 and $1$. The choice of convention that $f(0-)=0$ allows a sequence of functions that jump “right after time 0” to be possibly convergent in the M1’ topology. For example, one has $\mathrm{dist}(\mathbbm{1}_{[1/n,1]},\mathbbm{1}_{[0,1]})\to 0$ as $n\to\infty$, while the sequence $\mathbbm{1}_{[1/n,1]}$ is not convergent in the classical M1 topology, and in fact the M1’ topology is strictly weaker than the M1 topology. On the other hand, observe that $\mathbbm{1}_{[0,1/n]}$ is not convergent in $(\mathbb{D}[0,1],\mathrm{dist})$.
@@ 2.2.2. The $M$-oscillation
This theorem can be found in Chapter 12 of Whitt (2002), with the minor difference that our space of functions starts with an “initial jump” (recall that $f(0-)=0$ by our convention).
@@ 2.2.3. An exponential tightness criterion
For 1., we follow and adapt the approach of Billingsley (1968). Let $D_{n}=\{k2^{-n},0\leq k\leq 2^{n}\}$ be the dyadic numbers of level $n$. Then for $\beta>0$ and $\lambda>0$, by Markov’s inequality,
@@
which, by optimizing over $\lambda>0$ and taking a union bound, yields
@@
Setting $A_{n}=\max\{M(f(k2^{-n},(k+1)2^{-n},(k+2)2^{-n})),0\leq k\leq 2^{n}-2\}$, this shows that
$\mathbb{P}\left(A_{n}>2^{-n\beta}\right)\leq 2^{n}c\exp(-C^{\prime}(\alpha)\varepsilon^{-\alpha^{\prime}}2^{n(1-\alpha\beta)/(\alpha-1)}).$
Next, let $f\in\mathbb{D}[0,1]$ and, for $I\subset[0,1]$, let
@@ 3. Large deviations for the finite-dimensional marginal distributions
Let $\sigma=(t_{1},\ldots,t_{n})$, where $10<t_{1}<\dots<t_{n}<$ 1, be fixed. Under $\mathbb{P}$ the laws of $\varepsilon(\mathbbm{e}_{t_{1}},\dots,\mathbbm{e}_{t_{n}})$ satisfy an LDP in $\mathbb{R}^{n}$ with speed $\varepsilon^{-\alpha^{\prime}}$ and good rate function
@@ 3.1. Estimates for transition densities
First, (Sato, 1999, Equation (14.35)) entails that for every $x\geq 0$, we have
$c_{1}\exp\left(-c_{\alpha}x^{\alpha^{\prime}}\right)\leq p_{1}(-x)\leq c_{2}(1+x^{\frac{2-\alpha}{2\alpha-2}})\exp\left(-c_{\alpha}x^{\alpha^{\prime}}\right),$ (3.1)
and (Sato, 1999, Equation (14.34)) entails that
@@
Next, by (Monrad and Silverstein, 1979, Formula (3.20)), it holds that there exists a positive constant $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0})$ and for $0<a<b<+\infty$,
@@
These estimates will allow us to evaluate the densities involved in (2.3) in the large deviation regime. First we give an explicit formula for $p^{(0,\infty)}_{t}(x,y)$.
@@
Using the scaling properties of $p_{t}(x)$ and $q_{x}(t)$, we may deduce from Lemma 3.2 a bound on the error when we approximate $p^{(0,\infty)}_{t}(x,y)$ by $p_{t}(y-x)$, as follows.
@@
$\int_{0}^{t}q_{x}(s)p_{t-s}(y)\,\mathrm{d}s\leq\\ c_{2}\|p_{t/2}\|_{\infty}\int_{0}^{t/2}\frac{x\,\mathrm{d}s}{(t/2)^{\alpha}s^{1+1/\alpha}}(1+(x/s^{1/\alpha})^{\frac{2-\alpha}{2\alpha-2}})\exp\left(-c_{\alpha}(x/s^{1/\alpha})^{\alpha^{\prime}}\right)\\ +c_{2}\|p_{1}\|_{\infty}\frac{x}{(t/2)^{1+1/\alpha}}\left(1+\left(\frac{x}{(t/2)^{1/\alpha}}\right)^{\frac{2-\alpha}{2\alpha-2}}\right)\exp\left(-c_{\alpha}\left(\frac{x}{t^{1/\alpha}}\right)^{\alpha^{\prime}}\right)\int_{t/2}^{t}\frac{\,\mathrm{d}s}{(t-s)^{1/\alpha}}.$
@@
Proof of Proposition 3.1.
Since $J_{\sigma}$ is a good rate function on $\mathbb{R}^{n}$, (Serlet, 1997, Lemma 5) shows that it suffices to prove that, for every open subset $G\subset\mathbb{R}^{n}$,
@@
$\Psi_{\varepsilon}(x_{1},\dots,x_{n})=\prod_{i=1}^{n-1}p_{t_{i+1}-t_{i}}^{(0,\infty)}\Big{(}\frac{x_{i}}{\varepsilon},\frac{x_{i+1}}{\varepsilon}\Big{)}\times q_{x_{n}/\varepsilon}(1-t_{n}).$
@@
$p^{(0,\infty)}_{t_{i+1}-t_{i}}\left(\frac{x_{i}}{\varepsilon},\frac{x_{i+1}}{\varepsilon}\right)\\ \geq c_{1}\exp\left(-\frac{c_{\alpha}}{\varepsilon^{\alpha^{\prime}}}\left(\frac{(b_{i}-a_{i+1})^{\alpha}}{t_{i+1}-t_{i}}\right)^{\frac{1}{\alpha-1}}\right)-C(\eta)\exp\left(-\frac{c_{\alpha}}{\varepsilon^{\alpha^{\prime}}}(1-\eta)\left(\frac{a_{i}^{\alpha}}{t_{i+1}-t_{i}}\right)^{\frac{1}{\alpha-1}}\right),$
@@ 4. Exponential tightness for the normalized excursion
Now by (Monrad and Silverstein, 1979, (3.20)), we have that $\bar{J}_{1}(x)\sim\bar{c}x^{-\alpha}$ as $x\to\infty$ for some finite constant $\bar{c}$, and the desired uniform upper bound follows. ∎
@@
$\mathbb{P}\left(\sup_{1-\delta<t\leq 1}\varepsilon\mathbbm{e}_{t}\geq\gamma,\varepsilon\mathbbm{e}_{1-\delta}<\gamma\right)\\ \leq\alpha\Gamma\left(1-\frac{1}{\alpha}\right)\int_{0}^{\gamma/\varepsilon}\mathrm{d}x\,j_{1-\delta}(x)C(\eta,\gamma)\exp\bigg{(}-(1-\eta)c_{\alpha}\left(\frac{\gamma^{\alpha}}{\varepsilon^{\alpha}\delta}\right)^{\frac{1}{\alpha-1}}\bigg{)}.$
@@
Proof of Proposition 4.1.
Fix $\lambda>0$. By Lemma 4.3, for every $n\geq 1$, there exists a $\delta_{n}\in(0,1)$ such that for every $\varepsilon>0$ small enough
@@
For this choice of $\delta_{n}$, by Lemma 4.2 and Theorem 2.2, for every $n\geq 0$ there exists a compact set $K^{(n)}_{\lambda}$ of $\mathbb{D}[0,1]$ such that for every $\varepsilon>0$ small enough,
@@
5. Proof of Theorem 1.2
In this section, we prove Theorem 1.2. We will do this by combining the exponential tightness with a weaker form of the LDP, as we now explain. The weak topology $\mathcal{W}$ on $\mathbb{D}[0,1]$ is the topology generated by the basis of neighborhoods of the form
@@
where $f\in\mathbb{D}[0,1]$, $\varepsilon_{1},\ldots,\varepsilon_{k}>0$, and $t_{1},\ldots,t_{k}$ are elements of $[0,1]$ that are continuity points of $f$. Here, by convention, 00 is a continuity point of $f$ if and only if $f(0)=0$, which is consistent with our convention that $f(0-)=0$. Clearly this defines a Hausdorff topology, since two different elements of $\mathbb{D}[0,1]$ necessarily differ at some common continuity point. It is easy to see that a sequence $(f_{n},n\geq 0)$ that converges to a limit $f$ in $(\mathbb{D}[0,1],\mathrm{dist})$ also converges pointwise at every continuity point of $f$, and, therefore, will eventually belong to any given basic neighborhood $N(f,t_{1},\ldots,t_{k},\varepsilon_{1},\ldots,\varepsilon_{k})$ for the topology $(\mathbb{D}[0,1],\mathcal{W})$. Since open sets in the metric space $(\mathbb{D}[0,1],\mathrm{dist})$ can be characterized sequentially, this implies that the weak topology is coarser than the topology of $(\mathbb{D}[0,1],\mathrm{dist})$. Therefore, by (Dembo and Zeitouni, 1998, Corollary 4.2.6), and by the exponential tightness established in Proposition 4.1, Theorem 1.2 will follow from the following statement.
@@
The remainder of this section is thus devoted to the proof of this proposition, which follows the approach of Lynch and Sethuraman (1987) closely.
@@ 5.1. Facts about the rate function
where, by convention, $x_{n+1}=0$ and $t_{n+1}=1$. We let $J_{\sigma}(x_{1},\ldots,x_{n})=\infty$ if one of the $x_{i}$’s is negative. To ease notation, for $f:[0,1]\to\mathbb{R}$, we let $I_{\mathbbm{e}}^{\sigma}(f)=J_{\sigma}(f(t_{1}),\ldots,f(t_{n}))$. By the Dawson-Gärtner theorem (see (Dembo and Zeitouni, 1998, Theorem 4.6.1)), it follows from Proposition 3.1 that the laws of $(\varepsilon\mathbbm{e}_{t})_{t\in[0,1]}$ satisfy an LDP in $\mathbb{R}^{[0,1]}$ (with the product topology) as $\varepsilon\downarrow 0$, with speed $\varepsilon^{-\alpha^{\prime}}$ and good rate function
@@
We cannot immediately make use of this, since the domain of this rate function is not a space of càdlàg functions. However, let us prove some properties of the rate function $\widetilde{I}_{\mathbbm{e}}$ and, in particular, that its restriction to $\mathbb{D}[0,1]$ coincides with the rate function $I_{\mathbbm{e}}$ given in Theorem 1.2. To this end, we prove the following proposition.
@@
$M_{\mathbbm{e}}(f)=\sup_{\sigma\in\mathfrak{S}}I_{\mathbbm{e}}^{\sigma}(f)<\infty.$
@@
This statement should be compared with (Lynch and Sethuraman, 1987, Theorem 3.2), where the proof uses a martingale argument. We provide another elementary proof here, based on the Lebesgue differentiation theorem instead. For convenience, let $\Lambda(x)=c_{\alpha}(x_{-})^{\alpha^{\prime}}$ for all $x\in\mathbb{R}$.
@@
By the Lebesgue decomposition theorem, we may write $f_{\uparrow}=f_{\uparrow\mathrm{ac}}+f_{\uparrow\mathrm{sing}}$, where $f_{\uparrow\mathrm{ac}}$ is an absolutely continuous function and $f_{\uparrow\mathrm{sing}}$ is such that $\,\mathrm{d}f_{\uparrow\mathrm{sing}}$ is singular with respect to the Lebesgue measure. By the Lebesgue differentiation theorem, for Lebesgue-almost every $t\in[0,1]$ we have
@@
Since the weak topology is not first-countable, we must initially use nets to characterise the lower-semicontinuity of $I_{\mathbbm{e}}$. We first need to show that if $(f_{\lambda})$ is a net that converges to $f$ in the weak topology, with $f_{\lambda}\geq 0$ and $f_{\lambda}(1)=f(1)=0$ for every $\lambda$, then $\liminf_{\lambda}\ I_{\mathbbm{e}}(f_{\lambda})\geq I_{\mathbbm{e}}(f)$. Let $\sigma=(t_{1},\ldots,t_{k})$ be a subdivision of continuity points of $f$, so that $f_{\lambda}(t_{i})$ converges to $f(t_{i})$ for $1\leq i\leq k$, and $I^{\sigma}_{\mathbbm{e}}(f_{\lambda})=J_{\sigma}(f_{\lambda}(t_{1}),\ldots,f_{\lambda}(t_{k}))$ converges to $I^{\sigma}_{\mathbbm{e}}(f)=J_{\sigma}(f(t_{1}),\ldots,f(t_{k}))$. Since $I_{\mathbbm{e}}(f_{\lambda})\geq I^{\sigma}_{\mathbbm{e}}(f_{\lambda})$ by Proposition 5.2, this implies that $\liminf_{\lambda}I_{\mathbbm{e}}(f_{\lambda})\geq I^{\sigma}_{\mathbbm{e}}(f)$. Applying Proposition 5.2 once again allows us to conclude that $I_{\mathbbm{e}}$ is a rate function on $(\mathbb{D}[0,1],\mathcal{W})$, and therefore also on $(\mathbb{D}[0,1],\mathrm{dist})$.
@@
The proof follows that of (Lynch and Sethuraman, 1987, Theorem 3.5) closely. Since we know from Proposition 5.2 that $I^{\sigma}_{\mathbbm{e}}(A)\leq I_{\mathbbm{e}}(A)$ for every $\sigma\in\mathfrak{S}$ and every set $A$, let us assume, for a contradiction, that $\sup_{\sigma\in\mathfrak{S}}I^{\sigma}_{\mathbbm{e}}(F)<c<I_{\mathbbm{e}}(F)$ for some constant $c\in(0,\infty)$. For every subdivision $\sigma=(t_{1},\ldots,t_{k})\in\mathfrak{S}$, we may find some element $f_{\sigma}\in\mathbb{D}[0,1]$ such that $I^{\sigma}_{\mathbbm{e}}(f_{\sigma})<c$. Let $\hat{f}_{\sigma}$ be the piecewise affine interpolation of the values of $f_{\sigma}$ at times $10<t_{1}<\ldots<t_{k}<$ 1, with $\hat{f}_{\sigma}(0)=\hat{f}_{\sigma}(1)=0$ and of course $\hat{f}_{\sigma}(t_{i})=f_{\sigma}(t_{i}),1\leq i\leq k$. Then, plainly,
@@
We now have all the tools needed to prove Proposition 5.1. The proof is split into two lemmas which follow (Lynch and Sethuraman, 1987, Theorems 4.1 and 4.2) closely.